Write an equation of the line that contains the given point and has the given slope. Write the equation in slope intercept form.
1.(-6, -1), slope = -1
2.(-3, 2), slope = -3



Answer by Alan3354(69443)

Write an equation of the line that contains the given point and has the given slope. Write the equation in slope intercept form.
1.(-6,-1), slope = -1
------
y-y1 = m*(x-x1) where (x1,y1) is the pointi
y+1 = -1*(x+6)
y+1 = -x-6
Slope-intercept means solve for y.
y = -x-7
===================
2.(-3,2), slope = -3
Do it the same way as #1